### v2

@@ 15 Edinburgh (UK) 1983
“I believe that I reflect the consensus among statistical physicists when I assert a unique status for statistical physics among the varied subdisciplines of physics. That uniqueness derives from the generality of statistical physics; it is at the same time pervasive throughout physical theory, and absent a private, wholly-owned domain. [^79] These statements can be embraced even after almost fifty years. On the one hand, this attests to a positive peculiarity of statistical physics. On the other hand, this can lead to problems of recognition within the broader community of physicists. Organizationally we find our Commission to be the second oldest in IUPAP (and the first Commission dealing with a substantive subdiscipline), and we also find that statistical physics is the least intensively organized of all subdisciplines (witness the absence of a Statistical Physics Division within the American Physical Society). Ours is therefore an integrative role, unifying an international community of statistical physicists many, or even most, of whom count themselves simultaneously as members of one or another subdiscipline of physics. This interdisciplinary linkage has marked the field itself as well as its organizational structure, as some of the most significant contributions of statistical mechanics have transcended traditional boundaries. Perhaps the most recent and dramatic examples of this have been in the fields of broken symmetry and in the development of the techniques of the renormalization group."
@@ 8 Conclusion
However, while acknowledging the historical actors’ perspectives on their sub-disciplinary tradition, we have demonstrated the significant, albeit distinct, role played by the Brussels conference. This earlier gathering enabled members of the Commission at that time to engage in discussions that led to a reevaluation of the Commission’s role, culminating in a decision to rename and broaden its scope. Considering its pivotal role in setting the stage for the subsequent Florence conference, we suggest that the Brussels conference be recognized within the StatPhys series as StatPhys0. This approach aligns with practices adopted for labeling other international conferences, e.g. those on General Relativity and Gravitation. [^82] When the GR label was introduced, the 1957 Chapel Hill conference was retroactively designated GR1, while the 1955 Bern conference was identified as GR0, see [31]. Such a distinction would not diminish the Florence conference’s significant contributions to community building and field definition. Instead, it would offer a more comprehensive and nuanced account of Commission C3’s history.
@@
The community-building impact of the Florence conference was further magnified by the fact that, for many attendees, it was their first opportunity to convene since the devastation of World War II. The endeavor to move beyond the war’s atrocities and engage in a new communal experience with international peers was likely intensified by the event’s setting in Florence, a city celebrated as a jewel of the Renaissance. The organizers intentionally chose this location for its symbolic representation of the rejuvenation of international collaboration post-war, thereby amplifying the conference’s significance in the collective memory and fostering a spirit of renewal among the scientific community. [^83] All these aspects emerge clearly from the opening speeches given by Borghi [11], Kramers [30], and Polvani [47].
The conference held significant importance at the local level as well, serving as the inaugural step in a strategic endeavor by Italian physicists to revitalize Italian physics post-World War II through internationalization. The Florence conference was the first scientific event co-sponsored by IUPAP and the Italian Physical Society, marking the beginning of SIF’s forceful campaign to establish itself as a key collaborator with IUPAP in hosting international scientific conferences. This series of initiatives not only facilitated the promotion of Italian physics on the international stage but also played a crucial role in internationalizing its premier publication, Il Nuovo Cimento. Consequently, Il Nuovo Cimento ascended to prominence as a leading journal for particle physics in Europe in the subsequent years, reflecting the successful intertwining of local revitalization efforts with the broader objective of international engagement (see [32]).
@@
The success of the Florence conference significantly bolstered the Commission on Thermodynamics and Statistical Mechanics, setting a precedent for future IUPAP commissions. This pivotal event initiated a period of growth and self-identification for the Commission, as detailed in Section 7. The transformative years of the 1970s solidified the Commission’s identity, a foundation that has remained stable to the present day. Subsequent developments within the Commission can be viewed as natural progressions from these well-established roots. The focus on statistical mechanics, initiated at the Florence conference and further emphasized by the subsequent rebranding of the Commission’s general conferences as the StatPhys series, culminated in a significant renaming of the Commission itself in 1990 to the “C.3 Commission on Statistical Physics.” [^84] C.3 Commission on Thermodynamics and Statistical Mechanics, in Reports of International Commissions of IUPAP and the Inter-Union Committees presented to the 20th General Assembly, Dresden- September 1990, pp. 5-7, on p. 7; see also, IUPAP, General Report 1990, Appendix B, p. 15 , Series B2aa, Vol. 3, IuG. This modification was due to the understanding that, while statistical physics was initially related to the thermodynamics of fluids and magnets, from the 1970s the spectrum of research topics addressed within statistical physics became much broader, including “random media, dynamical systems, soft condensed matter, chaos, growth phenomena, and fractals.” As noted by the 1993 Commission C.3 report, “statistical physics has moved into more complex problems, which are closer to applications in diverse fields like biology, neural networks, computer science, astronomy or geophysics.”858585 C.3 Commission on Statistical Physics, in Reports of International Commissions of IUPAP and the Inter-Union Committees presented to the 21th General Assembly, Nara, Japan - September 1993, pp. 9-11, on p. 9, Series B2aa, Vol. 3, IuG. The most recent major change was introduced less than twenty years ago when IUPAP started financing a Young Scientist Prize for each commission with the same three-year periodicity of general conferences. [^86] For the list of awardees of Commission C3 Young Scientist Awards, see, https://archive2.iupap.org/commissions/c3-commission-on-statistical-physics/c3-awards/.
The oversight of StatPhys conference organization and the bestowment of Boltzmann and Young Scientist awards remain the primary functions of the C3 Commission to date, a role that has only been underscored by the Commission’s resolution to almost exclusively endorse type-A conferences. [^87] In the 2010s only a few type-B conferences were scheduled and StatPhys is the only type-A conference organized by C3. Private communication by Stefano Ruffo, former C3 chair, to one of the authors (PP) and information obtained from the C3 Commission’s website, https://iupap.org/who-we-are/internal-organization/commissions/c3-statistical-physics/.
@@
This strategic decision underscores the significance of StatPhys as the only conference championed by the C3 Commission for the whole community. The selection process for conference venues adheres to a rotational principle across continents, with the choice of country and city determined from submitted applications. There is no invitation procedure, with the decision being dictated by the desire to valorize an existing and growing local activities in statistical physics. In 2016, StatPhys26 in Lyon was the occasion to create a weekly newsletter in Statistical Physics, which has since informed the entire community about events and job offers. While this newsletter might seem a minor activity, it is still a direct offshoot of the StatPhys conferences, further demonstrating the long-lasting impact of a tradition inaugurated by the Florence conference.
@@ Acknowledgments
The text is partially based on research done by historians of science participating in the project “One Hundred Years of IUPAP: A History," coordinated by Jaume Navarro and one the authors (RL). We are very grateful for the support of IUPAP, which funded the digitization of its historical archive and the project’s workshops, which this paper has also benefited from. We are pleased to acknowledge the invaluable support of our colleagues Giovanni Battimelli, Adele La Rana, Giovanni Paoloni and Stefano Ruffo for their comments and suggestions. We are also very grateful to the following individuals who helped us in our research of archival and historical materials: Antonella Cotugno (Archives of the Department of Physics at Sapienza University), Francesca Gallori (Biblioteca Riccardiana, Florence), Leonardo Gariboldi (Archivio Polvani, University of Milan), Antonella Gasperini (Archivio Storico Osservatorio Arcetri), Erika Ghilardi (Archivio storico Foto Locchi), Alessia Glielmi (CNR Archives, Rome) Nicolien Karskens (Leiden University Libraries), Alessandra Lenzi (Museo Galileo, Florence), Anna Ranzi (Casa Museo Villa Monastero), Fioranna Salvadori (Archivio di Deposito e Storico, University of Florence), Maria Valdez (Documentary Information Services, Royal Netherlands Academy of Arts and Sciences), and Veronica Vestri (Deputazione di storia patria per la Toscana). Finally, we are very grateful to the President of the Italian Physical Society, Angela Bracco, and the society’s staff, especially Barbara Alzani, for the permission to access and cite the historical materials stored in the SIF offices in Bologna and their support during the research.
@@ Appendix A List of participants
This is the list of participants as reported by Polvani to Borghi on May 7, 1949. In a couple of cases (E. Bauer and C. Slater), expected keynote speakers didn’t give a lecture at the conference (or at least no account appears in the Proceedings).  [^88] See, Polvani to Borghi, May 7, 1949, 19/c-ASUFi. The presidents of the conference were H.A. Kramers and G. Polvani.
@@ E. Bauer (Paris)
I. Prigogine (Secretary, Bruxelles) Sur la perturbation de la distribution de Maxwell par des reactions chimiques
@@ A. Borsellino (Milano)
G. Careri (Roma)
N. Carrara (Firenze)
@@ B. De Finetti (Trieste)
S. Franchetti (Chairman, Firenze)
S. Gallone (Chairman, Milano)
Garikian (Bruxelles)
Gehenieau (Bruxelles)
M. Giusti (Secretary, Firenze)
L. van Hove (Bruxelles)
@@ F. London (Durhen, North Carolina)
G. Morpurgo (Chairman, Roma)
@@ F. Perrin (Paris)
P. Pinto (Milano)
L. Prinzi (Padova)
G. Puppi (Padova)
L. Rolla (Genova)
@@ E. Rudberg (Stockholm)
C. Salvetti (Chairman, Milano)
@@
G. Semerano (Padova)
P. Straneo (Genova)
@@ G. Todesco (Parma)
G. Toraldo di Francia (Chairman, Firenze)
P. Udeschini (Milano)
@@ G. Valle (Bologna)
G. Wataghin (Torino) Irreversible processes and the formation of nuclei
@@
G. Zin (Torino)
B. Zumino (Roma)

### v3

@@ 15 Edinburgh (UK) 1983
“I believe that I reflect the consensus among statistical physicists when I assert a unique status for statistical physics among the varied subdisciplines of physics. That uniqueness derives from the generality of statistical physics; it is at the same time pervasive throughout physical theory, and absent [as] a private, wholly-owned domain. [^92] These statements can be embraced even after almost fifty years. On the one hand, this attests to a positive peculiarity of statistical physics. On the other hand, this can lead to problems of recognition within the broader community of physicists. Organizationally we find our Commission to be the second oldest in IUPAP (and the first Commission dealing with a substantive subdiscipline), and we also find that statistical physics is the least intensively organized of all subdisciplines (witness the absence of a Statistical Physics Division within the American Physical Society). Ours is therefore an integrative role, unifying an international community of statistical physicists many, or even most, of whom count themselves simultaneously as members of one or another subdiscipline of physics. This interdisciplinary linkage has marked the field itself as well as its organizational structure, as some of the most significant contributions of statistical mechanics have transcended traditional boundaries. Perhaps the most recent and dramatic examples of this have been in the fields of broken symmetry and in the development of the techniques of the renormalization group."
@@ 8 Conclusion
However, while acknowledging the historical actors’ perspectives on their sub-disciplinary tradition, we have demonstrated the significant, albeit distinct, role played by the Brussels conference. This earlier gathering enabled members of the Commission at that time to engage in discussions that led to a reevaluation of the Commission’s role, culminating in a decision to rename it and broaden its scope. Considering its pivotal role in setting the stage for the subsequent Florence conference, we suggest that the Brussels conference be recognized within the StatPhys series as StatPhys0. This approach aligns with practices adopted for labeling other international conferences, e.g., those on General Relativity and Gravitation. [^95] When the GR label was introduced, the 1957 Chapel Hill conference was retroactively designated GR1, while the 1955 Bern conference was identified as GR0, see [47]. Such a distinction would not diminish the Florence conference’s significant contributions to community building and field definition. Instead, it would offer a more comprehensive and nuanced account of Commission C3’s history.
@@
The community-building impact of the Florence conference was further magnified by the fact that, for many attendees, it was their first opportunity to convene since the devastation of World War II. The endeavor to move beyond the war’s atrocities and engage in a new communal experience with international peers was likely intensified by the event’s setting in Florence, a city celebrated as a jewel of the Renaissance. The organizers intentionally chose this location for its symbolic representation of the rejuvenation of international collaboration post-war, thereby amplifying the conference’s significance in the collective memory and fostering a spirit of renewal among the scientific community. [^96] All these aspects emerge clearly from the opening speeches given by Borghi (1949), Kramers (1949), and Polvani (1949).
The conference held significant importance at the local level as well, serving as the inaugural step in a strategic endeavor by Italian physicists to revitalize Italian physics post-World War II through internationalization and to open up new research fields beyond cosmic rays as well as nuclear and subnuclear physics, the traditional workhorses of the Italian physics community. The Florence conference was the first scientific event co-sponsored by IUPAP and the Italian Physical Society, marking the beginning of SIF’s forceful campaign to establish itself as a key collaborator with IUPAP in hosting international scientific conferences. This series of initiatives not only facilitated the promotion of Italian physics on the international stage but also played a crucial role in internationalizing its premier publication, Il Nuovo Cimento. Consequently, Il Nuovo Cimento ascended to prominence as a leading journal for particle physics in Europe in the subsequent years, reflecting the successful intertwining of local revitalization efforts with the broader objective of international engagement [see 48].
@@
The success of the Florence conference significantly bolstered the Commission on Thermodynamics and Statistical Mechanics, setting a precedent for future IUPAP commissions. This pivotal event initiated a period of growth and self-identification for the Commission, as detailed in Section 7. The transformative years of the 1970s solidified the Commission’s identity, a foundation that has remained stable to the present day. Subsequent developments within the Commission can be viewed as natural progressions from these well-established roots. The focus on statistical mechanics, initiated at the Florence conference and further emphasized by the subsequent rebranding of the Commission’s general conferences as the StatPhys series, culminated in a significant renaming of the Commission itself in 1990 to the “C.3 Commission on Statistical Physics.” [^97] C.3 Commission on Thermodynamics and Statistical Mechanics, in Reports of International Commissions of IUPAP and the Inter-Union Committees presented to the 20th General Assembly, Dresden- September 1990, pp. 5-7, on p. 7; see also, IUPAP, General Report 1990, Appendix B, p. 15 , Series B2aa, Vol. 3, IuG. This modification was due to the understanding that, while statistical physics was initially related to the thermodynamics of fluids and magnets, from the 1970s the spectrum of research topics addressed within statistical physics became much broader, including “random media, dynamical systems, soft condensed matter, chaos, growth phenomena, and fractals.” As noted by the 1993 Commission C.3 report, “statistical physics has moved into more complex problems, which are closer to applications in diverse fields like biology, neural networks, computer science, astronomy or geophysics.”989898 C.3 Commission on Statistical Physics, in Reports of International Commissions of IUPAP and the Inter-Union Committees presented to the 21th General Assembly, Nara, Japan - September 1993, pp. 9-11, on p. 9, Series B2aa, Vol. 3, IuG. The most recent major change was introduced less than twenty years ago when IUPAP started financing a Young Scientist Prize for each commission with the same three-year periodicity of general conferences. [^99] For the list of awardees of Commission C3 Young Scientist Awards, see, https://archive2.iupap.org/commissions/c3-commission-on-statistical-physics/c3-awards/.
@@
The oversight of StatPhys conference organization and the bestowment of Boltzmann and Young Scientist awards remain the primary duties of the C3 Commission to date, a role that has only been accentuated by the Commission’s resolution to almost exclusively endorse type-A conferences.100100100In the 2010s only a few type-B conferences were scheduled and StatPhys is the only type-A conference organized by C3. Private communication by Stefano Ruffo, former C3 chair, to one of the authors (PP) and information obtained from the C3 Commission’s website, https://iupap.org/who-we-are/internal-organization/commissions/c3-statistical-physics/. This strategic decision underscores the significance of StatPhys as the only conference championed by the C3 Commission for the whole community. The selection process for conference venues adheres to a rotational principle across continents, with the choice of country and city determined from submitted applications. There is no invitation procedure, with the decision being dictated by the desire to valorize an existing and growing local activities in statistical physics. In 2016, StatPhys26 in Lyon was the occasion to create a weekly newsletter in Statistical Physics, which has since informed the entire community about events and job offers. While this newsletter might seem a minor activity, it is still a direct offshoot of the StatPhys conferences, further demonstrating the long-lasting impact of a tradition inaugurated by the Florence conference.
@@ Acknowledgments
The text is partially based on research done by historians of science participating in the project “One Hundred Years of IUPAP: A History," coordinated by Jaume Navarro and one the authors (RL). We are very grateful for the support of IUPAP, which funded the digitization of its historical archive and the project’s workshops, which this paper has also benefited from. We are pleased to acknowledge the invaluable support of our colleagues Giovanni Battimelli, Adele La Rana, Giovanni Paoloni and Stefano Ruffo for their comments and suggestions. We are also very grateful to the following individuals who helped us in our research of archival and historical materials: Antonella Cotugno (Archives of the Department of Physics at Sapienza University), Francesca Gallori (Biblioteca Riccardiana, Florence), Leonardo Gariboldi (Archivio Polvani, University of Milan), Antonella Gasperini (Archivio Storico Osservatorio Arcetri), Erika Ghilardi (Archivio storico Foto Locchi), Alessia Glielmi (CNR Archives, Rome) Nicolien Karskens (Leiden University Libraries), Alessandra Lenzi (Museo Galileo, Florence), Anna Ranzi (Casa Museo Villa Monastero), Fioranna Salvadori (Archivio di Deposito e Storico, University of Florence), Maria Valdez (Documentary Information Services, Royal Netherlands Academy of Arts and Sciences), and Veronica Vestri (Deputazione di storia patria per la Toscana). We want to express our deep gratitude to two anonymous referees for their deep reading of our paper and the insightful comments they provided. Finally, we are extremely grateful to the president of the Italian Physical Society, Angela Bracco, its honorary president, Luisa Cifarelli, and the society’s staff, especially Barbara Alcarani e Barbara Alzani, for the permission to access and cite the historical materials stored in the SIF offices in Bologna and their support during the research.
@@ Appendix A List of participants
This is the list of participants as reported by Polvani to Borghi on May 7, 1949. In a couple of cases (E. Bauer and C. Slater), expected keynote speakers didn’t give a lecture at the conference (or at least no account appears in the Proceedings). 101101101See, Polvani to Borghi, May 7, 1949, 19/c-ASUFi. The presidents of the conference were H.A. Kramers and G. Polvani.
@@ E. Bauer (Paris)
I. Prigogine (Secretary, Brussels) Sur la perturbation de la distribution de Maxwell par des reactions chimiques
@@ A. Borsellino (Milan)
G. Careri (Rome)
N. Carrara (Florence
@@ B. De Finetti (Trieste)
S. Franchetti (Chairman, Florence)
S. Gallone (Chairman, Milan)
Garikian (Brussels)
Gehenieau (Brussels)
M. Giusti (Secretary, Florence)
L. van Hove (Brussels)
@@ F. London (Durhen, North Carolina)
G. Morpurgo (Chairman, Rome)
@@ F. Perrin (Paris)
P. Pinto (Milan)
L. Prinzi (Padua)
G. Puppi (Padua)
L. Rolla (Genoa)
@@ E. Rudberg (Stockholm)
C. Salvetti (Chairman, Milan)
@@
G. Semerano (Padua)
P. Straneo (Genoa)
@@ G. Todesco (Parma)
G. Toraldo di Francia (Chairman, Florence)
P. Udeschini (Milan)
@@ G. Valle (Bologna)
G. Wataghin (Turin) Irreversible processes and the formation of nuclei
@@
G. Zin (Turin)
B. Zumino (Rome)